OpenStudy (anonymous):

what is the value of the expression -4x2 when x equals 2

OpenStudy (crabbyoldgamer):

Do you mean \[4x^2?\]

OpenStudy (anonymous):

ya

OpenStudy (crabbyoldgamer):

Substitute 2 in for the 4: \[-4x^2 \] when x = 2 becomes \[-4*(2)^2 = -4*4 = -16\]